# DrunkAgent: Stealthy Memory Corruption in LLM-Powered Recommender Agents

### v2

@@ 3.2 Generation Module
3.2.1 Adversarial Prompting Engineering
@@
This module designs the adversarial prompt for the surrogate recommendation agent to improve the quality of adversarial trigger generation. We craft prompt template, so that the surrogate knows ‘what needs to be done’ and ‘how to respond to’. As discussed in Section 2.2.1, it is challenging to obtain optimized memories of user agents and item agents in black-box attacks. To solve this and ensure the quality, we introduce general user memories and popular item memories.
Basic Prompt Template. The template $\widetilde{\mathcal{P}}$ is built with four components: Style Definition $\widetilde{p_{s}}$, Task Goal $\widetilde{p_{g}}$, Recommendation Instruction $\widetilde{p_{i}}$ and Format Constraint $\widetilde{p_{f}}$, as shown in the example of Fig. 2. The detailed descriptions are provided in Appendix B.1. Formally, $\widetilde{\mathcal{P}}=\widetilde{p_{s}}\oplus\widetilde{p_{g}}\oplus\widetilde{p_{i}}\oplus\widetilde{p_{f}}.$
@@
General User Memories. The triggers aim to be effective for general users, rather than niche user groups (e.g., those exclusively interested in Progressive Metal), to maximize the exposure rate of the target items. Hence, the user memory $\widetilde{m_{u}}$ is set with universal user descriptions, as the example shows in Fig. 2. This is also the initial memory of the user agent (zhang2024agentcf, ; huang2024foundation, ), which is accessible and aligns with the assumptions of target black-box RSs.
@@
Popular Item Memories. Besides the limited access of the optimized memories in practice, the inherent opacity of the victim systems renders the exact set of candidate items used for recommendations unknown. As the works (si2020shilling, ; liu2021adversarial, ; wang2025id, ) pointed out, popular items usually have high probabilities of appearing at the top of users’ recommendation lists. Intuitively, if target items can be ranked ahead of these popular items, they are more likely to be promoted to normal users. Therefore, we adopt popular items as candidate items to enhance the effectiveness of trigger optimizations. These popular items are publicly available, which are not beyond the black-box assumptions. Let $\mathcal{U}_{v}=\{u\in\mathcal{U}:y_{u,v}\neq 0\}$ indicates the set of users that have interacted with $v$ and $\mathcal{M}_{p}$ denotes a sequence of the features of popular items. Thus, $\mathcal{M}_{p}=\{m_{v}:v\in\mathcal{V}_{p}\}$, where $\mathcal{V}_{p}$ is a sequence of popular item IDs. The sequence is the corresponding subscript of the item popularity sequence (i.e., $\{|\mathcal{U}_{v}|\}_{v\in\mathcal{V}}$) in descending order. Raw item descriptions serve as a form of natural resource that can be leveraged to derive memories for popular items. However, such descriptions (ni2019justifying, ) are often too long that may exceed the context window limitations of the surrogate model and always contain too much redundant information that may increase the difficulty of subsequent optimizations. To tackle the issues, we propose to adopt a LLM (denoted as $f_{\text{LLM}_{\Psi}}(\cdot)$) to extract valuable item features from the raw descriptions and supplement the features based on its internal rich knowledge base to make them more informative to facilitate the optimizations. Let $\mathcal{C}_{p}$ represent the modified feature sequence of popular items, $\mathcal{C}_{p}=f_{\text{LLM}_{\Psi}}(\mathcal{X}_{p}\oplus\mathcal{M}_{p})$, where $\mathcal{X}_{p}$ denotes the prompts for feature acquisition. An example of producing a popular item memory is given in Fig. 2. Note that if the raw item descriptions are empty, we adopt other identity information including titles and categories.
@@
Agents always tend to select candidates positioned higher in the display list (zhang2024agentcf, ). To increase the trigger effectiveness, we place the other popular candidates before the target item, as shown in Fig. 2. In the previous sections, $m_{v}$ is commonly used to represent a set of features of item $v$. For better understanding the subsequent methods, for the target item $t$, we use different notations to distinguish between meta features and memories: $m_{t}$ denotes the metadata, while $\widetilde{m_{t}}$ denotes the memory (i.e., the adversarial trigger). As a result, the adversarial prompt $\widetilde{\mathcal{X}_{t}}$ can be formulated as
$\widetilde{\mathcal{X}_{t}}=\widetilde{\mathcal{P}}\oplus\widetilde{m_{u}}\oplus\mathcal{C}_{p}\oplus m_{t}.$ (2)
3.2.2 Greedy Search Algorithm
The algorithm is used to optimize the adversarial description triggers, inspired by the recent successes in jailbreak attacks on general-purpose LLMs (zou2023universal, ; liu2023autodan, ). Despite this inspiration, DrunkAgent fundamentally differs in key aspects such as attack goals and victim models. Additional details are given in Appendix A. To fill the gaps, the algorithm incorporates task-aligned initialization and refines the search space tailored to manipulate the behaviors of the target agents.
@@ 3.2.2 Greedy Search Algorithm
~$\widetilde{\mathcal{M}}_{c}=f_{\text{LLM}_{\Psi}}(\widetilde{\mathcal{X}_{c}}\oplus m_{t}),$ (3)
where ~$\widetilde{\mathcal{M}}_{c}=\{\widetilde{m}_{t_{i}}\}$ represents a sequence of trigger candidates obtained from the LLM.
@@
$\mathcal{L}_{\widetilde{m}_{t_{i}}^{(k-1)}}=-\log\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{R}_{t}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)})),$ (4)
@@
Feature Integration Stage. The algorithm samples a subset of |~$|\widetilde{\mathcal{M}}_{c}|-n$ input candidates by conducting probability-based random selection. The candidates with higher scores will have a greater probability of being selected, where the selection probability of each candidate is calculated via a softmax function, i.e., $\text{softmax}(\widetilde{m}_{t_{i}}^{(k-1)})=\frac{e^{s_{i}^{(k-1)}}}{\sum_{j=1}^{|\widetilde{\mathcal{M}}_{c}|}e^{s_{j}^{(k-1)}}}$. For the obtained subset, the algorithm splits the candidates at random positions into a list of text slices and performs the random exchange of the slices between pairwise texts, allowing features from different candidates to be fused, as shown in Fig. 2. Specifically, this fusion creates new combinations of words, phrases, and sentences, enhancing the feature diversity of candidates, and hence contributing to search for an optimal trigger.
@@
$\widetilde{m_{t}}^{*}=\mathop{\arg\max}_{\widetilde{m}_{t_{i}}^{\mathcal{E}}}(\{s_{i}^{\mathcal{E}}\}),$ (5)
@@ 3.3 Strategy Module
$\widetilde{q}=\widetilde{q_{f}}\oplus\widetilde{q_{c}}\oplus\widetilde{q_{g}}\oplus\widetilde{q_{n}}\oplus\widetilde{q_{s}},$ (6)
where an example of $\widetilde{q}$ is given in Fig. 2.
3.3.2 Adversarial Strategy Optimization.
The arrangement of the defined strategies needs to be optimized to improve the attack effectiveness. DrunkAgent obtains the optimal strategy $\widetilde{q}^{*}$ by maximizing the expected probability that the agent when influenced by adversarial modifications, performs a malicious action for a given input query,
$\mathbbm{E}_{\widetilde{q}\sim\pi_{\widetilde{q}}}[\ \mathbbm{I}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{X}_{s}}\oplus\widetilde{q}\oplus\widetilde{m_{t}}^{*})=\widetilde{a_{t}})],$ (7)
where $\pi_{\widetilde{q}}$ denotes the distribution of adversarial strategies, $\mathbbm{I}(\cdot)$ is an indicator function and $\widetilde{\mathcal{X}_{s}}$ is an adversarial prompt for simulating memory updates of item agents during the agents’ environmental interactions. $\widetilde{a_{t}}$ is the desired malicious action for the injected strategy $\widetilde{q}$, i.e., the target item agent is unable to fulfill specific roles and is incapable of perceiving, learning and self-evolving from the interaction environment (e.g., fail in language-based signal back-propagation (zhang2024agentcf, )). An example of $\widetilde{q}^{*}$ is given in Fig. 2. To sum up, the adversarial description of the target item is $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$. The overall optimization procedure of DrunkAgent will be provided in Algorithm 1.
The above design further reveals how memory mechanisms in agent-based recommender systems can be influenced under black-box conditions. These insights may support future efforts toward developing memory-aware defenses (e.g., using customized deep neural network detectors (wu2020densely, ; yang2021dualnet, ; yang2021hunter, ; yang2021deep, )) and enhancing the robustness of agent-powered RSs chua2024ai .
4 Experiments
4.1 Experimental Setup
Datasets Selection. To comprehensively evaluate DrunkAgent, we use three datasets that are text-intensive containing real-world data and widely-adopted in the current recommendation studies (huang2024foundation, ; zhang2024agentcf, ). They are CDs & Vinyl, Office Products and Musical Instruments of Amazon Review Data (ni2019justifying, ). Their statistics are given in Appendix C.1.1. We additionally evaluate our method on Yelp dataset [^2] https://www.yelp.com/dataset to demonstrate its generalizability across domains in the appendix. The target items are randomly sampled from the item pool on each dataset. The universality of attacks across different target items is further demonstrated in Appendix C.2.1.
Baseline Attack Methods. Since we are the first work to gain insight into the robustness of the generative agent-powered RSs, there is no any baselines. To comprehensively evaluate the effectiveness of DrunkAgent, we follow the state-of-the-art attack on LLM-based RSs (zhang2024stealthy, ) and compare DrunkAgent with the benign status and six black-box attack methods, where the perturbations range from character-level: DeepwordBug (gao2018black, ) and PuncAttack (formento2023using, ), word-level: TextFooler (jin2020bert, ) and BertAttack (li2020bert, ) to sentence-level: TrivialInsertion and ChatGPTAttack (zhang2024stealthy, ). To maintain a fair comparison, the methods perform perturbations on the descriptions of the target items and are optimized by the surrogate model that is the same as DrunkAgent. More detailed implementations are provided in Appendix C.1.2.
Targeted Recommender Systems. We adopt three different agentic paradigms (zhang2024agentcf, ) as victim RSs to evaluate the effectiveness of DrunkAgent across diverse agent designs and recommendation tasks: AgentCF, a standard CF paradigm; AgentRAG, which augments the agents with retrieval; and AgentSEQ, which captures temporal dynamics in user behaviors for sequential recommendations. Given that the victim RSs are totally different from as well as more powerful and context-rich than the surrogate, a strict black-box setting is maintained for sufficient evaluations. To ensure a fair evaluation, we follow the default implementations, as elaborated in Appendix C.1.3.
The configuration of DrunkAgent is given in Appendix C.1.4. We evaluate its performance using widely-used HR@$\mathcal{K}\uparrow$, NDCG@$\mathcal{K}\uparrow$, and sentence perplexity score $\downarrow$. The specific definitions of the metrics are in Appendix C.1.5. More experiments and demonstrations (e.g., ablation studies and the sensitivity of critical parameters) can be found in the appendix.
Table 1: Attack Transferability. HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ of different attacks against various black-box victim LLM-powered agent-based RSs on real-world datasets. We use bold fonts to denote the best performance. The attacks with higher HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ have excellent transferability.
@@ 4.1 Experimental Setup
4.2 Attack Transferability

### v3

@@ 3.2. Generation Module
3.2.1. Recommendation Context Construction
@@
This module introduces adversarial prompts for the surrogate recommendation agent to improve the quality of adversarial trigger generation. Specifically, we craft a template to establish a recommendation-style interaction context. As discussed in Section 2.2.1, it is challenging to obtain optimized memories of user agents and item agents within the black-box attacks. Therefore, we propose to incorporate general user memories and popular item memories to approximate those memories to ensure both effectiveness and reliability.
Basic Prompt Template. The template $\widetilde{\mathcal{P}}$ is built with four components: 1) Style Definition $\widetilde{p_{s}}$, which introduces role information into the context to elicit domain-specific responses; 2) Task Goal $\widetilde{p_{g}}$, which specifies the ranking objective, aligning the task with recommendation (Zhang et al., 2019); 3) Recommendation Instruction $\widetilde{p_{i}}$, which provides explicit guidance for comparing user preferences with candidate items; and 4) Format Constraint $\widetilde{p_{f}}$, which restricts the output format to reduce undesirable outputs and facilitate subsequent optimization. An illustration is given in Fig. 2. Formally, $\widetilde{\mathcal{P}}=\widetilde{p_{s}}\oplus\widetilde{p_{g}}\oplus\widetilde{p_{i}}\oplus\widetilde{p_{f}}$.
@@
General User Memories. The triggers aim to be effective for general users, rather than niche user groups (e.g., those exclusively interested in Progressive Metal), to maximize the exposure rate of the target items. Hence, the user memory $\widetilde{m_{u}}$ is set with universal user descriptions, as the example shows in Fig. 2. This is also the initial memory of the user agent (Zhang et al., 2024b; Huang et al., 2024), which is accessible and aligns with the assumptions of target black-box RSs.
Popular Item Memories. Besides the limited access of the optimized memories in practice, the inherent opacity of the victim systems renders the exact set of candidate items used for recommendations unknown. As the works (Si and Li, 2020; Liu and Larson, 2021; Wang et al., 2025) pointed out, popular items usually have high probabilities of appearing at the top of users’ recommendation lists. Intuitively, if target items can be ranked ahead of these popular items, they are more likely to be promoted to normal users. Therefore, we adopt popular items as candidate items to enhance the effectiveness of trigger optimizations. These popular items are publicly available, which are not beyond the black-box assumptions.
@@
Let $\mathcal{U}_{v}=\{u\in\mathcal{U}:y_{u,v}\neq 0\}$ indicates the set of users that have interacted with $v$ and $\mathcal{M}_{p}$ denotes a sequence of the features of popular items. Thus, $\mathcal{M}_{p}=\{m_{v}:v\in\mathcal{V}_{p}\}$, where $\mathcal{V}_{p}$ is a sequence of popular item IDs. The sequence is the corresponding subscript of the item popularity sequence (i.e., $\{|\mathcal{U}_{v}|\}_{v\in\mathcal{V}}$) in descending order. Raw item descriptions serve as a form of natural resource that can be leveraged to derive memories for popular items. However, such descriptions (Ni et al., 2019) are often too long that may exceed the context window limitations of the surrogate model and always contain too much redundant information that may increase the difficulty of subsequent optimizations. To tackle the issues, we propose to adopt a LLM (denoted as $f_{\text{LLM}_{\Psi}}(\cdot)$) to extract valuable item features from the raw descriptions and supplement the features based on its internal rich knowledge base to make them more informative to facilitate the optimizations. Let $\mathcal{C}_{p}$ represent the modified feature sequence of popular items, $\mathcal{C}_{p}=f_{\text{LLM}_{\Psi}}(\mathcal{X}_{p}\oplus\mathcal{M}_{p})$, where $\mathcal{X}_{p}$ denotes the prompts for feature refinement. An example of producing a popular item memory is given in Fig. 2.
@@
Agents always tend to select candidates positioned higher in the display list (Zhang et al., 2024b). To increase the trigger effectiveness, we place the other popular candidates before the target item, as shown in Fig. 2. In the previous sections, $m_{v}$ is commonly used to represent a set of features of item $v$. For better understanding the subsequent methods, for the target item $t$, we use different notations to distinguish between meta features and memories: $m_{t}$ denotes the metadata, while $\widetilde{m_{t}}$ denotes the memory (i.e., the adversarial trigger). As a result, the adversarial prompt $\widetilde{\mathcal{X}_{t}}$ can be formulated as
(2) $\widetilde{\mathcal{X}_{t}}=\widetilde{\mathcal{P}}\oplus\widetilde{m_{u}}\oplus\mathcal{C}_{p}\oplus m_{t}.$
3.2.2. Greedy Search Algorithm
We introduce a greedy search algorithm to optimize the adversarial description triggers. Starting from attack goal–aligned seeds, the algorithm iteratively selects, recombines, and polishes candidates, progressively refining the search space to produce highly effective textual triggers for promoting target items.
@@ 3.2.2. Greedy Search Algorithm
(3) $\widetilde{\mathcal{M}}_{c}=f_{\text{LLM}_{\Psi}}(\widetilde{\mathcal{X}_{c}}\oplus m_{t}),$
where $\widetilde{\mathcal{M}}_{c}=\{\widetilde{m}_{t_{i}}\}$ represents a sequence of trigger candidates.
@@
(4) $\mathcal{L}_{\widetilde{m}_{t_{i}}^{(k-1)}}=-\log\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{R}_{t}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)})),$
@@
Feature Integration Stage. The algorithm samples a subset of $|\widetilde{\mathcal{M}}_{c}|-n$ input candidates by conducting probability-based random selection. The candidates with higher scores will have a greater probability of being selected, where the selection probability of each candidate is calculated via a softmax function, i.e., $\text{softmax}(\widetilde{m}_{t_{i}}^{(k-1)})=\frac{e^{s_{i}^{(k-1)}}}{\sum_{j=1}^{|\widetilde{\mathcal{M}}_{c}|}e^{s_{j}^{(k-1)}}}$. For the obtained subset, the algorithm splits the candidates at random positions (e.g., at punctuation marks) into a list of text slices and performs the random exchange of the slices between pairwise texts, as shown in Fig. 2, allowing features from different candidates to be fused. Specifically, this fusion creates new combinations of words, phrases, and sentences, enhancing the feature diversity of candidates, and hence contributing to search for an optimal trigger.
Algorithm 1 Optimization Procedure of DrunkAgent
1: Obtain the adversarial prompt $\widetilde{\mathcal{X}_{t}}$ with Eq. (2)
@@
4:  Quality Estimation Stage:
5:  Calculate $s_{i}^{(k-1)}$ for each input trigger $i$ with Eq. (4)
6:  Maintain top-$n$ candidates in descending order of scores
7:  Feature Integration Stage:
8:  Sample a subset of $|\widetilde{\mathcal{M}}_{c}|-n$ input candidates via Softmax
9:  Randomly exchange of the slices between pairwise texts
10:  Linguistic Enrichment Stage:
11:  Optimize and polish the $|\widetilde{\mathcal{M}}_{c}|-n$ previous combinations via LLMs
12:  Obtain $|\widetilde{\mathcal{M}}_{c}|$ triggers for the next epoch optimization
13:end while
15: Define the adversarial strategy $\widetilde{q}$ with Eq. (6)
18:end while
19: Obtain the optimal strategy $\widetilde{q}^{*}$ when Eq. (7) is maximum
Linguistic Enrichment Stage. Given that interleaving two candidates can result in the combinations that are awkward or lack fluency, a LLM is introduced to polish the combinations. In this way, not only the original semantic meanings are preserved, but also the clarity, coherence and natural flow of the text are improved, thereby enhancing the concealment of the attack. Moreover, the LLM expands the candidates’ corpus to avoid the problem of language-based vanishing gradient causing the algorithm to fall into local optimum. Since excessive use of new data may lead to a language-based gradient explosion that makes the optimization unstable, the length limitation is introduced into the prompt $\widetilde{\mathcal{X}_{r}}$. Formally, for each new candidate $\widetilde{m}_{t_{i}}^{(k)}$, $\widetilde{m}_{t_{i}}^{(k)}=f_{\text{LLM}_{\Psi}}(\widetilde{\mathcal{X}_{r}}\oplus\widetilde{m}_{t_{i}}^{(k-1)^{\prime}})$, where $\widetilde{m}_{t_{i}}^{(k-1)^{\prime}}$ denotes the resulting combination from the previous stage. Consequently, $n$ well-performing trigger candidates from the first stage and the subset $\{\widetilde{m}_{t_{i}}^{(k)}:i\ \text{is the index of the second stage's candidate}\}$ with the length $|\widetilde{\mathcal{M}}_{c}|-n$ will be optimized in the next epoch $k+1$.
@@
(5) $\widetilde{m_{t}}^{*}=\mathop{\arg\max}_{\widetilde{m}_{t_{i}}^{\mathcal{E}}}(\{s_{i}^{\mathcal{E}}\}),$
@@ 3.3. Strategy Module
3.3.1. Perturbation Strategy Definition.
[…]
(6) $\widetilde{q}=\widetilde{q_{f}}\oplus\widetilde{q_{c}}\oplus\widetilde{q_{g}}\oplus\widetilde{q_{n}}\oplus\widetilde{q_{s}}$,
where an example of $\widetilde{q}$ is given in Fig. 2.
3.3.2. Adversarial Strategy Optimization.
The organization of the defined strategies needs to be optimized to improve the effectiveness of ‘get the target agent drunk’. DrunkAgent obtains the optimal strategy $\widetilde{q}^{*}$ by maximizing the expected probability that the agent when influenced by adversarial modifications, performs a malicious action for a given input query,
[…]
where $\pi_{\widetilde{q}}$ denotes the distribution of adversarial strategies, $\mathbbm{I}(\cdot)$ is an indicator function and $\widetilde{\mathcal{X}_{s}}$ is an adversarial prompt template for simulating memory updates of item agents during the agents’ environmental interactions. $\widetilde{a_{t}}$ is the desired malicious action for the injected strategy $\widetilde{q}$, i.e., the target item agent is unable to fulfill specific roles and is incapable of perceiving, learning and self-evolving from the interaction environment (e.g., fail in language-based signal back-propagation (Zhang et al., 2024b)). An example of $\widetilde{q}^{*}$ is given in Fig. 2.
To sum up, the adversarial description of the target item is $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$. For transparency and reproducibility, the overall optimization procedure of DrunkAgent is given in Algorithm 1.
The above design further reveals how memory mechanisms in agent-based recommender systems can be influenced under black-box conditions. These insights may support future efforts toward developing memory-aware defenses (e.g., using customized deep neural network detectors (Wu et al., 2020; Yang et al., 2021b, a; Yang, 2021)) and enhancing the robustness of agent-powered RSs (Chua et al., 2024).
4. Experiments
4.1. Experimental Setup
4.1.1. Datasets Selection.
To comprehensively evaluate DrunkAgent, we use three datasets that are text-intensive containing real-world data and widely-adopted in the current recommendation studies (Huang et al., 2024; Zhang et al., 2024b). They are CDs & Vinyl, Office Products and Musical Instruments of Amazon Review Data (Ni et al., 2019). We additionally evaluate our method on Yelp dataset [^2] https://www.yelp.com/dataset to demonstrate its generalizability across domains in the appendix. The target items are randomly sampled from the item pool on each dataset.
4.1.2. Baseline Attack Methods.
Since we are the first work to gain insight into the robustness of the generative agent-powered RSs, there is no any baselines. To comprehensively evaluate the effectiveness of DrunkAgent, we follow the state-of-the-art attack on LLM-based RSs (Zhang et al., 2024c) and compare DrunkAgent with the benign status and six black-box attack methods, where the perturbations range from character-level: DeepwordBug (Gao et al., 2018) and PuncAttack (Formento et al., 2023), word-level: TextFooler (Jin et al., 2020) and BertAttack (Li et al., 2020) to sentence-level: TrivialInsertion and ChatGPTAttack (Zhang et al., 2024c). To maintain a fair comparison, the methods perform perturbations on the descriptions of the target items and are optimized by the surrogate model that is the same as DrunkAgent.
4.1.3. Targeted Recommender Systems.
We adopt three different agentic paradigms (Zhang et al., 2024b) as victim RSs to evaluate the effectiveness of DrunkAgent across diverse agent designs and recommendation tasks: AgentCF, a standard CF paradigm; AgentRAG, which augments the agents with retrieval; and AgentSEQ, which captures temporal dynamics in user behaviors for sequential recommendations. Given that the victim RSs are totally different from as well as more powerful and context-rich than the surrogate, a strict black-box setting is maintained for sufficient evaluations. To ensure a fair evaluation, we follow the default implementations.
4.1.4. Evaluation Metrics
[…]
4.1.5. Configuration of DrunkAgent
For the surrogate $f_{\text{Agent}_{\Phi}}(\cdot)$ backbone, we adopt the open-source model Meta-Llama-3-8B-Instruct [^3] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct, a recent and well-established LLM, which provides a favorable trade-off between efficiency and quality. In contrast, OpenAI’s gpt-4-turbo [^4] https://platform.openai.com/docs/models/gpt-4-turbo is adopted to implement $f_{\text{LLM}_{\Psi}}(\cdot)$, offering stronger effectiveness while maintaining reasonable cost. We set $\mathcal{E}=20$, $|\widetilde{\mathcal{M}}_{c}|=10$ and $n=5$. All random seeds in the evaluation are set to 2024 and the victim LLMs’ temperature is set to 0 for reproducibility. For all the other parameters, we keep the default values.
More implementation details, experiments and demonstrations (e.g., ablation studies and the sensitivity of critical parameters) can be found in the appendix.
4.2. Attack Transferability
4.2.1. Overall Transferability.